# Mirage Sources and Large TeV Halo-Pulsar Offsets: Exploring the Parameter Space

### v2

@@ II Method Description
To characterise the asymmetric propagation of particles, we conduct test-particle simulations by calculating the trajectories of 100 TeV electrons, the radiation spectrum of which typically peaks at approximately 10 TeV via the inverse Compton (IC) scattering off the cosmic microwave background (CMB). We employ a first-principle approach by injecting electrons into a turbulent magnetic field and determining their motion based on the Lorentz force [24, 25, 26]. We employ GPU-acceleration methods, allowing the code to operate in a highly parallelized manner. To preserve GPU on-board memory, we calculate the magnetic field anew at every time step for every particles’ position. While this approach is time-intensive, it offers a greater accuracy compared to the method which specifies fields on a discrete spatial lattice in advance and interpolates while computing. Here, we thoroughly explain in detail the process of conducting the simulation in terms of particle motion and field generation.
@@ II.2 Simulating Particle Propagations
We employ the original Boris pusher which avoids the phase error [see e.g., Boris A pusher in 27] to preserve phase space volume. Since the electric field in the turbulent plasma may be neglected [24], the Boris pusher can be written as
$\bm{x}^{\prime}:$ $=\bm{x}_{n-1}+\frac{1}{2}\bm{v}_{n-1}\cdot\Delta t$ (1)
$\bm{t}:$ $=\tan\frac{\theta}{2}\hat{\bm{b}}$ (2)
$\gamma\bm{v}^{\prime}:$ $=\gamma\bm{v}_{n-1}+\gamma\bm{v}_{n-1}\times\bm{t}$ (3)
$\gamma\bm{v}_{n}$ $=\gamma\bm{v}^{\prime}+\frac{2}{1+|\bm{t}|^{2}}\left(\gamma\bm{v}^{\prime}\times\bm{t}\right)$ (4)
$\bm{x}_{n}$ $=\bm{x}^{\prime}+\frac{1}{2}\bm{v}_{n}\cdot\Delta t$ (5)
where $\hat{\bm{b}}:=\bm{B}/|\bm{B}|$ is the unit vector, $\gamma$ the electron Lorenrz factor, $n$ the time step, $\bm{x}^{\prime}$, $\bm{t}$, $\bm{v}^{\prime}$ are three intermediate variables. A time step of 1/100 gyration period is utilized in the simulation. We also take synchrotron/IC losses into consideration. The synchrotron loss of electrons with Lorentz factor $\gamma$ reads
@@ III.1 Methods
Figure 1: Variation of the source numbers $\xi$ with different plasma parameters: (a) dependence on the safety factor $q$ or inverse aspect ratio $\alpha$; (b) scaling with the toroidal magnetic field $B_{t}$; (c) variation with the connection length $L_{c}$; (d) relationship with the ratio of turbulent to neoclassical transport. The results demonstrate how $\xi$ depends on different parameters and provide insights into the characteristic scale of turbulence structures in the plasma.

### v4

@@ II Method Description
To characterise the asymmetric propagation of particles, we conduct test-particle simulations by calculating the trajectories of 100 TeV electrons, the radiation spectrum of which typically peaks at approximately 10 TeV via the inverse Compton (IC) scattering off the cosmic microwave background (CMB). We employ a first-principle approach by injecting electrons into a turbulent magnetic field and determining their motion based on the Lorentz force [16, 12, 17]. We employ GPU-acceleration methods, allowing the code to operate in a highly parallelized manner. Here, we thoroughly explain in detail the process of conducting the simulation in terms of particle motion and field generation.
@@ II.2 Simulating Particle Propagations
We employ the original Boris pusher which avoids the phase error [see e.g., Boris A pusher in 32] to preserve phase space volume. Since the electric field in the turbulent plasma may be neglected [16], the Boris pusher can be written as
$\boldsymbol{x}^{\prime}:$ $=\boldsymbol{x}_{n-1}+\frac{1}{2}\boldsymbol{v}_{n-1}\cdot\Delta t$ (1)
$\boldsymbol{t}:$ $=\tan\frac{\theta}{2}\hat{\boldsymbol{b}}$ (2)
$\gamma\boldsymbol{v}^{\prime}:$ $=\gamma\boldsymbol{v}_{n-1}+\gamma\boldsymbol{v}_{n-1}\times\boldsymbol{t}$ (3)
$\gamma\boldsymbol{v}_{n}$ $=\gamma\boldsymbol{v}^{\prime}+\frac{2}{1+|\boldsymbol{t}|^{2}}\left(\gamma\boldsymbol{v}^{\prime}\times\boldsymbol{t}\right)$ (4)
$\boldsymbol{x}_{n}$ $=\boldsymbol{x}^{\prime}+\frac{1}{2}\boldsymbol{v}_{n}\cdot\Delta t$ (5)
where $\hat{\boldsymbol{b}}:=\boldsymbol{B}/|\boldsymbol{B}|$ is the unit vector, $\gamma$ the electron Lorenrz factor, $n$ the time step, $\boldsymbol{x}^{\prime}$, $\boldsymbol{t}$, $\boldsymbol{v}^{\prime}$ are three intermediate variables. A time step of 1/100 gyration period is utilized in the simulation. We also take synchrotron/IC losses into consideration. The synchrotron loss of electrons with Lorentz factor $\gamma$ reads
@@ III.1 Methods
Figure 1: Variation of the source numbers $\xi$ with different plasma parameters: (a) electron injection index $\alpha$, (b) total strength of the magnetic field, (c) coherence length of the turbulent field and (d) regular-to-turbulent ratio. The results demonstrate how $\xi$ depends on different parameters and provide insights into the characteristic scale of turbulence structures in the plasma.